Can someone help me solve this expression. 25a^2 - 23a - 2
Answer by checkley77(12844)

25a^2-23a-2
POSSIBLE FACTORS of 25 =1&25, -1&-25, 5&5, -5&-5
POSSIBLE FACTORS of -2 =1&-2, 2&-1
(25a-1)(a+2)=(25a^2-49a-2) not a solution.
(5a-1)(5a+2)=(25a^2+5a-2) ditto.
(-25a+1)(-a-2)=(25a^2+49a-2) ditto.
(-5a+1)(-5a-2)=25a^2+5a-2) ditto
(25a-2)(a+1)=(25a^2-23a-2) solution.